A compound containing only carbon and hydrogen is found to contain 4.1315 g of carbon and 0.8685 g of hydrogen. Determine the empirical formula of hydrocarbon
1
Expert's answer
2018-02-10T06:18:07-0500
CxHy
w(C) = 4.1315 g / (4.1315 g + 0.8685 g) = 0.8263
w(H) = 0.8685 g / (4.1315 g + 0.8685 g) = 0.1737
x : y = 0.8263/12,0107 : 0.1737/1,00794 = 0.06877 : 0.17233 = 4 : 10
C4H10